For each molecule give the absorbance for each important functional group.1. For each molecule give the absorbance for each important FG (a) (b) CH-C CH он (d) (c) H NH2 (c) Amphetamine A synthetic c
Answer #1

Answer (9) c=0 - 1705-1715 cm 2c-4 = 2900 - 3000 cml Sp? 21 3=C-H = 3300 cm- CH₃-CECH { 2 CEC = 2200 cm ñ C-H = 2000 - 3000cm